Fermi National Laboratory
 

Operation Summary for March 5, 2010 - March 8, 2010

 

The Integrated Luminosity for the period from 3/1/10 to 3/8/10 was 61.86 pb-1.
NuMI reported receiving 8.33E18 protons on target during this same period.

Friday March 5

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (15.9E10), with the Antiproton Source stacking (42.8E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, Operations established store 7651 at 8:11 AM with an Average Initial Luminosity (AIL) of 315E30. Operations received a Booster enclosure FIRUS alarm at 8:11 AM and prepared the Booster for access. Booster personnel assisted the Fire Department and found an arcing high voltage cable for a RF station (BRF4). Booster personnel repaired the cable and Operations reestablished Booster beam by 11:00 AM. TeV personnel moved the D0 interaction point by a small amount in the afternoon.

During the evening shift, Operations disabled the MTest beam at 6:02 PM. Operations terminated store 7651 at 8 PM. Operations put the TeV into HEP shot setup but it was halted when a wet engine failed at sector A2 and required an 8-hour rebuild.

Saturday March 6

The midnight shift began with the TeV off for wet engine work at A2, with Operations monitoring the stash (351.2E10), with Pbar stacking (14.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, A2 wet engine repairs continued. Cryo system recovery was underway by 5:30 AM with Operations establishing a TeV ramp by 7 AM. Operators conducted orbit checks until near to the end of the shift. Antiproton transfers to the Recycler were completed when appropriate; the stash limit was reached by 3:15 AM. Operations established beam to MTest at 4:45 AM, Operations put the TeV through a dry squeeze at 7:23 AM.

During the day shift, after the orbits checks Operations put the TeV into shot setup at 9:12 AM. Operations established store 7655 at 10:43 AM with an AIL of 347E30 and then established beam to MTest. The only sources of minor downtime were due to various RF trips in Linac and Booster.

During the evening shift, a Linac RF station (LRF5) went to zero with no indication as to why; operators ran it back up.

Sunday March 7

The midnight shift began with Operations monitoring store 7655 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations disabled the MTest beam at 12:07 AM. Operations terminated store 7655 at 1 AM and put the TeV into shot setup. Operations established store 7656 at 2:30 AM with an AIL of 319E30. Operations established beam to MTest at 4:37 AM.

During the day shift, an MTest quadrupole power supply (MT4Q2) wouldn’t turn on. A power supply person came in and replaced a ground fault fuse.

During the evening shift, Operations disabled the MTest beam at 6:15 PM. Operations terminated store 7656 at 7 PM and then put the TeV into shot setup, but a network communications problem at C4 halted the setup. Operators called in controls personnel. Controls personnel replaced the network switch at C4 and the TeV was ready for beam operations by 9:30 PM. Operations put the TeV through a dry squeeze and then into shot setup at 10 PM. Operations established store 7658 at 11:15 PM with an AIL of 318E30. CDF personnel reported that their detector experienced a power supply failure during the antiproton load; they will require a 3-hour access for repairs in between stores.

Monday March 8

The midnight shift began with Operations monitoring store 7658 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:20 AM.

The Plans for Monday

The plans for today is to terminate store 7658 at 9 AM and prepare the TeV A-E, CDF, and D0 enclosures for access. After all repairs have been made, the accelerator complex will resume stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/08/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory